# Linking emitted drops to collective bursting bubbles across a wide range of bubble size distributions

### v1

@@ 4.1 Direct attribution between measured bubbles and drops
The lack of large drops for two of the narrow-banded cases (peaking at radii of of 2 mm and 3 mm) can be explained by either missing drops due to the small field of view, missing drops that fall back to the surface before reaching the measurement region (Néel and Deike,, 2022), or by a physical collective effect where neighbors may somehow limit production efficiency in these nearly monodisperse conditions.
@@
To take a step in investigating a potential collective effect that may modulate the bubble to drop size link or affect overall production efficiencies, we look at the broad-banded cases and compare cases with 32 versus 16 bubbling needles (with the same flowrate per needle, so effectively halving the injected volume of bubbles). Aside from decreasing the injected volume, decreasing the number of needles allows the bubbles to arrive at the surface spaced apart from each other, so that most bubbles burst individually instead of clustered, as visualized in figure 1(f). Looking at the size distributions for these two cases, shown in the bottom row of figure 6 (g-i), we observe that a smaller number of bursting bubbles (g) results in a drop size distribution that is very similar to that of the clustered case, especially for the submicron drops (h). Looking at the efficiencies (i), there is a consistent factor of 1.5–2 between the 32 and 16 needle cases across all bubble sizes, indicating that a collective effect could potentially be limiting the drop production efficiency.
@@
Through the attribution of measured drops to measured bubbles analyzed in this section, we have observed that the broad banded cases show an overall agreement between our calculated bursting efficiencies and those found from the single bubble scaling laws for film flap and jet drop production, with a possible reduction of efficiency for the collective bursting (g-i). The narrow-banded cases demonstrate that submicron drops are produced by bubbles with radii smaller than $\sim$1 mm (d-f), and that nearly monodisperse clusters may display a reduced efficiency in jetting for bubbles with radii above $\sim$100 µm and an increased efficiency for the smaller bubbles (c, f).
@@ 4.2 Using individual bubble scaling laws to predict drops produced by measured bubbles
To predict the drops produced by the collection of bursting bubbles, we consider the approach from Lhuissier and Villermaux, (2012), initially proposed for production of film drops via a centrifuge mechanism and extended to jet drops (Berny et al.,, 2021) and film flap drops (Deike et al.,, 2022). For a specific drop production process, the total drop size distribution $N_{d}(r_{d})$ produced by an ensemble of bubbles is obtained by integrating individual bursting scalings over the size distribution of bursting bubbles $N_{b}(R_{b})$ in the size range [$R_{b1}$, $R_{b2}$]:
@@
where $N_{b}(R_{b})$ is the measured bubble bursting size distribution, $\langle r_{d}\rangle(R_{b})$ is the mean drop size produced by a bursting bubble of size $R_{b}$, $n_{d}(R_{b})$ is the number of drop produced by a bubble bursting of size $R_{b}$, and $p(r_{d}/\langle r_{d}\rangle,R_{b})$ is the probability density function of drop sizes produced by a bubble of size $R_{b}$, assumed to be a Gamma distribution of known order (Lhuissier and Villermaux,, 2012; Berny et al.,, 2021; Deike et al.,, 2022).
We perform the integration, using the measured bubble size distribution, for each individual drop production mechanism, as summarized in table 3. Scalings for the number and size of drops produced by bursting of single bubbles have been developed for three main modes of drop production: film centrifuge, film flap, and jet drop production. The integration for each mechanism over many bubble sizes is similar to the analysis performed in Deike et al., (2022) (using assumed/modeled bubble size distributions) and Néel et al., (2022) (measuring only larger scales of drops and bubbles), as well as for individual mechanisms in Lhuissier and Villermaux, (2012) (film centrifuge drops, assumed/modeled bubble size distribution) Blanco–Rodríguez and Gordillo, (2020) (jet drops), and Gañán-Calvo, (2023) (extended jet drops using assumed/modeled bubble size distributions).
@@
Mechanism [$R_{b1}$, $R_{b2}]$ (µm) $\langle r_{d}\rangle(R_{b})$ (µm) $n_{d}(R_{b})$ $\Gamma$ order
film flap [70, 1000] $0.40\ \left(\frac{R_{b}}{l_{c}}\right)^{1/3}$ 40 6
@@
jet (B-R) [30, 2300] $0.20\ R_{b}\left(1-\left({\frac{R_{b}}{0.033^{2}l_{\mu}}}\right)^{-1/4}\right)$ $30\left(\frac{R_{b}}{l_{\mu}}\right)^{-1/3}$ 11
@@
Figure 7: Measured bubble ($N_{b}(R_{b})$) and drop ($N_{d}(r_{d})$) size distributions for the 2 mm injected bubble cases with (a) broad-banded and (b) narrow-banded distributions. Overlaid colored lines represent the predicted drop size distribution from each production mechanisms for individual bubble bursting, integrated over many bubble sizes using the measured bubble size distribution. Bubble size ranges used in the integration for each mechanism are labeled in the same color on the bubble size distributions.
@@
We start with jet drop production, where drops are created by fragmentation of the jet formed by the focusing of capillary waves into the collapsed bubble cavity. For bubbles of sizes between 30 µm and the capillary length ($l_{c}$ $\sim$ 2.3 mm for the artificial seawater solution), which are able to produce jet drops (Berny et al.,, 2020; Walls et al.,, 2015), we apply the formulation developed by Gañán-Calvo, (2017) for the resulting drop size and Berny et al., (2021) for number of drops (coherent with experimental data from Ghabache et al., (2014); Ghabache and Séon, (2016); Spiel, 1997b ) with a modified prefactor. The modified prefactor means that bubbles of radii $R_{b}$ = [30, 2300] µm are predicted to produce $n_{d}$ = [3, 0.6] jet drops per bubble, which is lower than the $n_{d}$ = [11, 3] drops from the Berny et al., (2021) fit to various data for individual bubble bursting, but is in agreement with the collective bursting experiment from Wang et al., (2017) (see figure 1 from their paper).
@@
For the broad-banded case, shown in the top row of figure 7 (a; bubbles on the left, drops on the right), the jet drop scaling laws (blue lines) match the measured drop size distribution (in black) very well for drops of radii approximately 1 µm to 500 µm. Because the magnitude of the distributions lines up well using the 0.6–3 jet drops produced per bubble across the radius range (where smaller bubbles produce more drops), the average number of jet drops produced by many collective bursting events is likely lower than the 3–11 drops that a single bubble is physically able to produce (across the same radius range). Each bursting event is not necessarily as efficient as every other event. In the collective bursting setup, some bubbles may be disturbed by the rising or bursting of their neighbors or by water fluctuations that could affect the cavity collapse and jet formation, which would reduce the overall drop production efficiency. The jet drop size formulation developed in Blanco–Rodríguez and Gordillo, (2020), based on capillary wave selection, was also tested (blue dotted line). Since it yields very similar results (as the drop sizes as a function of bubble sizes were derived using the same experimental and numerical data), we use the equation from Gañán-Calvo, (2017) for the rest of the analysis.
@@
Next we consider the mode of film drop production proposed in Jiang et al., (2022), where they report submicron drops produced by bursting bubbles in the approximate radius range $R_{b}$ = [70, 1000] µm, which they attribute to a film flapping mechanism. As shown in figure 7, the size scaling and order of magnitude associated with the film flap drops (orange) match well with the measured distribution (black) for the broad-banded case.
@@
Finally, for film drop production through a centrifuge mechanism, Lhuissier and Villermaux, (2012) describe the fragmentation of ligaments formed as the cap of a bubble with radius $R_{b}>0.4l_{c}$ retracts. We observe that the predicted distribution (green) does not fit the data well, underpredicting the distribution for the broad-banded case. The film centrifuge scaling leads to significantly fewer drops than the jet drop counterpart in the same size range due to the fact that there are fewer larger bubbles (above the capillary length) where the centrifuge mechanism is effective. The film centrifuge mechanism is also not likely to be the dominant mechanism for the production of smaller drops.
@@
Looking now at the narrow-banded case presented in figure 6(b), we observe that the film flap results again match the peak and magnitude relatively well (orange line), only slightly below the experimental data (black line). The jet scalings (blue) approximate the correct drop magnitude at their peak, around 1 µm, but overpredict larger drops (which again might be related to limits in measuring larger drops in the experiment). The film centrifuge mechanism again does not describe the data, this time overpredicting the number of drops. This significant overprediction may be related to fact that the centrifuge mechanism can frequently eject drops horizontally, reducing their measured efficiency as these drops likely fall back to the water without being transported upwards into the measurement region (Néel and Deike,, 2022; Shaw and Deike,, 2024).
@@
Figure 8: Bursting bubble ($N_{b}(R_{b})$) and drop ($N_{d}(r_{d})$) size distributions for the cases: (a) Broad-3 mm, (b) Broad-half volume, (c) Narrow-3 mm, and (d) Narrow-250 µm, with representative surface images inlaid for each case. The predicted drop size distributions from individual bubble bursting scaling laws are plotted in color for film flapping (orange) and jet (blue) drop production, with the bounds of the associated portions of the bubble size distribution shown with dashed lines of the same color.
@@ 4.3 Jet and film drop individual bubble bursting scaling laws for all cases
For the cases with narrower distributions, shown in rows (c, d) of the same figure, the measured and predicted distributions match well in the size range of film flap drops but are less coherent for the region described by jet drop production. In the large bubble case (row (c)), again fewer large jet drops are measured than predicted. For the small bubbles, the prediction overestimates drops around radius $r_{d}$ = 10 µm. As discussed in section 4.1 for the direct attribution, the narrow-banded cases reveal that the assumed size link between bubbles and drops is not entirely accurate in this collective bursting setup when large aggregate of bubbles of nearly the same size are present. Indeed, it seems that the predicted jet drop distribution for the cases (c, d) may be missing drops above radius $r_{d}$ = 1 µm, which can be seen from the gap between the predicted and measured distributions where the film and jet curves overlap. Néel and Deike, (2021) and Néel et al., (2022) observed a similar shift/broadening in the peak size of ejected drops for collective bursting of millimetric bubbles in surfactant solutions of various concentrations, which remains unexplained. These inconsistencies may be explained by a surfactant or collective effect and require further investigation. We note also that the predicted jet drop production is sensitive to the chosen critical Laplace number ($La^{*}$, the drop ejection threshold), shifting the predicted drop size distribution to somewhat smaller sizes if this value is increased.
@@ 5 Conclusion
Overall, we have studied the link between bulk bubbles, surface bubbles, and drops in a collective bursting setup, and we have experimentally confirmed the applicability of mechanistic ocean spray emissions functions based on individual bursting mechanisms (Deike,, 2022; Deike et al.,, 2022). We have leveraged many combined measurements and cases of diverse bubble size distributions and underwater conditions to analyze the attribution of drops to bursting bubbles. The potential collective effect suggested by these experiments will be studied further through dynamical measurements of bursting in bubble rafts. Future work using the same setup will explore how the link between bursting bubbles and emitted drops changes for solutions of different surfactant and temperatures, further extending the applicability of physics-based sea spray emissions functions (Deike et al.,, 2022) to open ocean conditions in presence of biological activity (Burrows et al.,, 2014) and taking steps to understand sea spray emissions by bubble bursting in other increasingly realistic configurations for ocean conditions.
@@
Appendix A Number of drops and bubbles for all cases
@@
The different conditions explored throughout the paper are characterized by very different numbers of bubbles at the surface, and therefore very different numbers of emitted drops, as illustrated by figure 9, showing the number of drops as a function of the number of bubbles with radii smaller than 2 mm. As indicated in the figure and already mentioned, our experiments span a range of number of bubbles over four orders of magnitude. An averaged drop production efficiency for each surface bubble configuration can be evaluated by an approximate linear relationship between the number of drops and the number of bubbles, as indicated by the solid line in figure 9. This linear fit suggests similar averaged efficiencies among all cases, with about 40 drops per bursting bubble. We also observe that the two cases of broad-banded bubble size distributions with significant clustering (cases Broad-2 mm and Broad-3 mm) show a reduced efficiency (pink and black symbols being below the linear trend).
@@
Figure 9: Number of bubbles with radii below 2 mm ($n_{R_{b}<2mm}$) versus total number of drops ($n_{d}$), both per unit volume, for each case explored in the paper. The gray line represents $n_{d}=40~{}n_{R_{b}<2mm}$, where the coefficient represents an average drop production efficiency, assuming drop production by bubbles with $R_{b}<$ 2 mm.

### v2

@@ 4.1 Direct attribution between measured bubbles and drops
The lack of large drops for two of the narrow-banded cases (peaking at radii of 2 mm and 3 mm) can be explained by either missing drops due to the small field of view or by a physical collective effect where neighbors may somehow limit production efficiency in these nearly monodisperse conditions.
@@
To take a step in investigating a potential collective effect that may modulate the bubble to drop size link or affect overall production efficiencies, we look at the broad-banded cases and compare cases with 32 versus 16 bubbling needles (“Broad-2 mm” and “Broad-2 mm-half volume” – the cases have the same flowrate per needle, so we effectively halve the injected volume of bubbles). Aside from decreasing the injected volume, decreasing the number of needles allows the bubbles to arrive at the surface spaced apart from each other, so that most bubbles burst individually instead of clustered, as visualized in figure 1(f). Looking at the size distributions for these two cases, shown in the bottom row of figure 6 (g-i), we observe that a smaller number of bursting bubbles (g) results in a drop size distribution that is very similar to that of the clustered case, especially for the submicron drops (h). Looking at the efficiencies (i), the 16 needle case has a higher efficiency than the 32 needle case across all bubble sizes, indicating that a collective effect could potentially be limiting the drop production efficiency.
@@
Through the attribution of measured drops to measured bubbles analyzed in this section, we have observed that the broad-banded cases show an overall agreement between our calculated bursting efficiencies and those found from the single bubble scaling laws for film flap and jet drop production, with a possible reduction of efficiency for the collective bursting (g-i). The narrow-banded cases demonstrate that submicron drops are produced by bubbles with radii smaller than $\sim$1 mm (d-f), and that nearly monodisperse clusters may display a reduced efficiency in jetting for bubbles with radii above $\sim$100 µm and an increased efficiency for the smaller bubbles (c, f).
@@ 4.2 Using individual bubble scaling laws to predict drops produced by measured bubbles
To predict the drops produced by the collection of bursting bubbles, we consider the approach from Lhuissier & Villermaux (2012), initially proposed for production of film drops via a centrifuge mechanism and extended to jet drops (Berny et al., 2021) and film flap drops (Deike et al., 2022). For a specific drop production process, the total drop size distribution $N_{d}(r_{d})$ produced by an ensemble of bubbles is obtained by integrating individual bursting scalings over the size distribution of bursting bubbles $N_{b}(R_{b})$ in the size range [$R_{b1}$, $R_{b2}$]:
@@
where $N_{b}(R_{b})$ is the measured bubble bursting size distribution, $\langle r_{d}\rangle(R_{b})$ is the mean drop size produced by a bursting bubble of size $R_{b}$, $n_{d}(R_{b})$ is the number of drop produced by a bubble bursting of size $R_{b}$, and $p(r_{d}/\langle r_{d}\rangle,R_{b})$ is the probability density function of drop sizes produced by a bubble of size $R_{b}$, assumed to be a Gamma distribution of known order (Lhuissier & Villermaux, 2012; Berny et al., 2021; Deike et al., 2022). Sensitivity of the resulting drop size distributions to the order of the Gamma distribution and to the chosen radius bounds ($R_{b1}$ and $R_{b2}$) are discussed in appendix B.
We perform the integration, using the measured bubble size distribution, for each individual drop production mechanism, as summarized in table 3. Scalings for the number and size of drops produced by bursting of single bubbles have been developed for three main modes of drop production: film centrifuge, film flap, and jet drop production. The integration for each mechanism over many bubble sizes is similar to the analysis performed in Deike et al. (2022) (using assumed/modeled bubble size distributions) and Néel et al. (2022) (measuring only larger scales of drops and bubbles), as well as for individual mechanisms in Lhuissier & Villermaux (2012) (film centrifuge drops, assumed/modeled bubble size distribution) Blanco–Rodríguez & Gordillo (2020) (jet drops), and Gañán-Calvo (2023) (extended jet drops using assumed/modeled bubble size distributions).
@@
Mechanism [$R_{b1}$, $R_{b2}]$ (µm) $\langle r_{d}\rangle(R_{b})$ (µm) $n_{d}(R_{b})$ Order of $\Gamma$
distribution
film flap [70, 1000] $0.40\ \left(\frac{R_{b}}{l_{c}}\right)^{1/3}$ 40 4
@@
jet (B-R) [30, 500] $0.20\ R_{b}\left(1-\left({\frac{0.033^{2}R_{b}}{l_{\mu}}}\right)^{-1/4}\right)$ $30\left(\frac{R_{b}}{l_{\mu}}\right)^{-1/3}$ 11
@@
Figure 7: Measured bubble ($N_{b}(R_{b})$) and liquid drop ($N_{d}(r_{d})$) size distributions for the 2 mm injected bubble cases with (a) broad-banded and (b) narrow-banded distributions. Overlaid colored lines represent the predicted drop size distribution from each production mechanism for individual bubble bursting (see details in table 3), integrated over many bubble sizes using the measured bubble size distribution. Bubble size ranges used in the integration for each mechanism are labeled in the same color on the measured bubble size distributions. Note that the jet (B-R) formulation is only applied in the range $R_{b}$ = [30,500] µm.
@@
We start with jet drop production, where drops are created by fragmentation of the jet formed by the focusing of capillary waves into the collapsed bubble cavity. For bubbles of sizes between 30 µm and the capillary length ($l_{c}$ $\sim$ 2.3 mm for the artificial seawater solution), which are able to produce jet drops (Berny et al., 2020; Walls et al., 2015), we apply the theoretical formulation developed by Gañán-Calvo (2017) for the resulting drop size. Because there is no theoretical prediction for the number of jet drops produced by the bursting bubble, we consider the scaling proposed by Berny et al. (2021) for the number of drops (coherent with experimental data from Ghabache et al. (2014); Ghabache & Séon (2016); Spiel (1997b)) with a modified prefactor. The modified prefactor means that bubbles of radii $R_{b}$ = [30, 2300] µm are predicted to produce $n_{d}$ = [3, 0.6] jet drops per bubble, which is lower than the $n_{d}$ = [11, 3] drops from the Berny et al. (2021) fit to various data for individual bubble bursting, but is in agreement with the collective bursting experiment from Wang et al. (2017) (see figure 1 from their paper).
@@
For the broad-banded case, shown in the top row of figure 7 (a; bubbles on the left, drops on the right), the jet drop scaling laws (blue lines) match the measured drop size distribution (in black) very well for drops of radii approximately 1 µm to 500 µm. Because the magnitude of the distributions lines up well using the 0.6–3 jet drops produced per bubble across the radius range (where smaller bubbles produce more drops), the average number of jet drops produced by many collective bursting events is likely lower than the 3–11 drops that a single bubble is physically able to produce (across the same radius range). Each bursting event is not necessarily as efficient as every other event. In the collective bursting setup, some bubbles may be disturbed by the rising or bursting of their neighbors or by water fluctuations that could affect the cavity collapse and jet formation, which would reduce the overall drop production efficiency. The jet drop size formulation developed in Blanco–Rodríguez & Gordillo (2020), based on capillary wave selection, was also tested (blue dotted line). Since it yields very similar results over the range of bubble radii we measure where the scaling is valid (as the drop sizes as a function of bubble sizes were derived using the same experimental and numerical data in both formulations), we use the equation from Gañán-Calvo (2017) for the rest of the analysis.
@@
Next we consider the mode of film drop production proposed in Jiang et al. (2022), where they report submicron drops produced by bursting bubbles in the approximate radius range $R_{b}$ = [70, 1000] µm, which they attribute to a film flapping mechanism. As shown in figure 7, the size scaling and order of magnitude associated with the film flap drops (orange) match well with the measured distribution (black) for the broad-banded case.
@@
Finally, for film drop production through a centrifuge mechanism, Lhuissier & Villermaux (2012) describe the fragmentation of ligaments formed as the cap of a bubble with radius $R_{b}>0.4l_{c}$ retracts. We observe that the predicted distribution (green) does not describe the data well, falling below the measured distribution for the broad-banded case. The film centrifuge scaling leads to significantly fewer drops than the jet drop counterpart (in the size range where they both mechanisms are applicable) due to the fact that there are fewer larger bubbles (above the capillary length) where the centrifuge mechanism is effective. Note that the film centrifuge scaling does not predict submicron drops.
@@
Looking now at the narrow-banded case presented in figure 6(b), we observe that the film flap results again match the peak and magnitude relatively well (orange line), only slightly below the experimental data (black line). The jet scalings (blue) approximate the correct drop magnitude at their peak, around 1 µm, but overpredict some larger drops. While this overprediction could be related to limits in measuring larger drops in the experiment, we also note that if a truncated Gamma distribution is used, such that $p(r_{d}/\langle r_{d}\rangle,R_{b})$ = 0 for bubbles of radii $R_{b}>$ 2 mm, the resulting predicted drop size distribution describes the large drops well (shown in appendix B for the Broad-2 mm and Narrow-2 mm cases). The film centrifuge mechanism again does not describe the data, this time overpredicting the number of drops. This significant overprediction may be related to fact that the centrifuge mechanism can frequently eject drops horizontally, reducing their measured efficiency as these drops likely fall back to the water without being transported upwards into the measurement region (Néel & Deike, 2022; Shaw & Deike, 2024).
@@
Figure 8: Bursting bubble ($N_{b}(R_{b})$) and liquid drop ($N_{d}(r_{d})$) size distributions for the cases: (a) Broad-3 mm, (b) Broad-2 mm-half volume, (c) Narrow-3 mm, and (d) Narrow-250 µm, with representative surface images inlaid for each case. The predicted drop size distributions from individual bubble bursting scaling laws are plotted in color for film flapping (orange) and jet (blue) drop production (see details in table 3), with the bounds of the associated portions of the bubble size distribution shown with dashed lines of the same color.
@@ 4.3 Jet and film drop individual bubble bursting scaling laws for all cases
For the cases with narrower distributions, shown in rows (c, d) of the same figure, the measured and predicted distributions match well in the size range of film flap drops but are less coherent for the region described by jet drop production. In the large bubble case (row (c)), again fewer large jet drops are measured than predicted. For the small bubbles, the prediction overestimates drops around radius $r_{d}$ = 10 µm. As discussed in section 4.1 for the direct attribution, the narrow-banded cases reveal that the assumed size link between bubbles and drops is not entirely accurate in this collective bursting setup when large aggregate of bubbles of nearly the same size are present. Indeed, it seems that the predicted jet drop distribution for the cases (c, d) may be missing some drops below radius $r_{d}$ = 1 µm, which can be seen from the gap between the predicted and measured distributions where the film and jet curves overlap. Néel & Deike (2021) and Néel et al. (2022) observed a similar shift/broadening in the peak size of ejected drops for collective bursting of millimetric bubbles in surfactant solutions of various concentrations, which remains unexplained. These inconsistencies may be explained by a surfactant or collective effect and require further investigation. We note also that the predicted jet drop production is sensitive to the chosen critical Laplace number ($La^{*}$, the drop ejection threshold), shifting the predicted drop size distribution to somewhat smaller sizes if this value is increased.
@@ 5 Conclusion
Overall, we have studied the link between bulk bubbles, surface bubbles, and drops in a collective bursting setup, and we have experimentally confirmed the applicability of mechanistic ocean spray emissions functions based on individual bursting mechanisms (Deike, 2022; Deike et al., 2022). We have leveraged many combined measurements and cases of diverse bubble size distributions and underwater conditions to analyze the attribution of drops to bursting bubbles. The potential collective effect suggested by these experiments will be studied further through dynamical measurements of bursting in bubble rafts. Future work using the same setup will explore how the link between bursting bubbles and emitted drops changes for solutions of different surfactant and temperatures, further extending the applicability of physics-based sea spray emissions functions (Deike et al., 2022) to open ocean conditions in presence of biological activity (Burrows et al., 2014) and taking steps to understand sea spray emissions by bubble bursting in other increasingly realistic configurations for ocean conditions.
@@
Data availability statement: Data used to prepare this work is publicly available at https://doi.org/10.34770/7s70-0v84.
@@
Appendix A Average drop production efficiencies
@@
The different conditions explored throughout the paper are characterized by very different numbers of bubbles at the surface, and therefore very different numbers of emitted drops, as illustrated by figure 9, showing the number of drops as a function of the number of bubbles with radii smaller than 2 mm. As indicated in the figure and already mentioned, our experiments span a range of number of bubbles around three orders of magnitude. An averaged drop production efficiency for each surface bubble configuration can be evaluated by an approximate linear relationship between the number of drops and the number of bubbles, as indicated by the solid line in figure 9. This linear fit suggests similar average efficiencies among all cases, with about 40 drops per bursting bubble. We also observe that the two cases of broad-banded bubble size distributions with significant clustering (cases Broad-2 mm and Broad-3 mm) show a reduced efficiency (pink and black symbols being below the linear trend).
Experiments were also performed for various air flowrates of 200 cm3/min, 400 cm3/min, and 800 cm3/min for the bubbler configuration, and all cases yield drop production efficiencies similar to those reported in figure 9, with no significant measurable effect of the air flowrate on the production of small drops in our setup.
@@
Figure 9: Total number of drops ($n_{d}$) versus the number of bubbles with radii below 2 mm ($n_{R_{b}<2mm}$), both per unit volume, for each case explored in the paper. The gray line represents $n_{d}=40~{}n_{R_{b}<2mm}$, where the coefficient represents an average drop production efficiency, assuming drop production by bubbles with $R_{b}<$ 2 mm.
Appendix B Uncertainty in drop production efficiencies and sensitivity of predicted drop size distributions
When performing the direct attribution of drops to bubbles from the data (section 4.1) and calculating drop production efficiency for film and jet drops, there are uncertainties in the measurement of the bubble size distribution and drop size distribution as well as uncertainties related to the chosen radius bounds of integration. As discussed in the experimental methods, the experimental size distributions are statistically converged, and the statistics are highly reproducible when repeating the same experiment. From measurements of bubble and drop size distributions, the estimated uncertainty on the efficiency values shown in figure 6 is around a factor 2.
We note, however, that the predicted drop size distributions shown in sections 4.2 and 4.3 (or the figure 6 efficiencies calculated in a particular size bucket) can be quite sensitive to the choice of bubble and/or drop radius ranges that are integrated for each drop production mechanism (film or jet). To calculate the efficiencies presented in figure 6, the total ranges of assumed jet drops and corresponding bubbles are split into approximately equal log-spaced buckets, of proportional sizes between the drop and bubbles. Shrinking/expanding/shifting these buckets will cause the efficiency values to shift up or down slightly, which could change the physical interpretation. The interpretation we propose results from our understanding of individual bursting processes discussed in the literature and from the consistency between the message of the direct attribution (section 4.1) and the drop size distributions predicted by the individual bubble bursting scaling laws combined with the measured bubble size distributions (sections 4.2, 4.3). We note that the overall average drop production efficiencies (see appendix A) are relatively constant across all cases. This overall efficiency is dominated by small drop production by bubbles of intermediate size, which is shown in both our methods of drop attribution to bursting bubbles for all cases, and gives confidence in our chosen interpretation.
Regarding the analysis presented in sections 4.2 and 4.3, we provide further details on the sensitivity of the drop size distributions predicted using the single bubble scalings (used to generate the colored curves presented in figures 7 and 8), shown in figure 10. We test the sensitivity to (a) the order of the Gamma distribution representing $p(r_{d}/\langle r_{d}\rangle,R_{b})$ and to (b) the lower and (c) upper radius bounds chosen for the film flapping and jet drop production mechanisms. From these figures, we observe that the order of the Gamma distribution primarily affects the shape of the curve at the small drop-size cutoff, and the chosen upper radius bound shifts the large drop-size cutoff. The predicted distributions are much more sensitive to the chosen lower radius bound for each production mechanism, which more significantly affects the magnitude of the predicted distributions for the film flap drops and is critical in defining the location of the peak for small jet drops (and therefore also the important overlap between film flap and jet scalings).
We note that we use the Gamma distribution to represent the statistics of drops emitted by a single bubble bursting, but another type of distribution, such as a lognormal distribution, could be used. As discussed in Berny et al. (2021), when integrating over a wide range of bubbles sizes, the scaling for the mean size and number of drops emitted dominates over the type of function used, and the type of function will mostly modulate the result near the bounds of integration. The Gamma distribution has the practical advantage of allowing for analytical integration when considering bubble size distributions represented as power laws, as discussed in Lhuissier & Villermaux (2012).
Figure 10: Bursting bubble ($N_{b}(R_{b})$) and drop ($N_{d}(r_{d})$) size distributions for the Broad-2 mm (top row) and Narrow-2 mm (bottom row) cases. The predicted drop size distributions from individual bubble bursting scaling laws are plotted in color for film flapping (orange) and jet (blue, from G-C scaling) drop production. The sensitivity of the predicted drop size distributions to parameters of the model/integration (equation 2) is shown for: (a) the order of the Gamma distribution $p(r_{d}/\langle r_{d}\rangle,R_{b})$ (varied between 2 and 11), (b) the lower bound, $R_{b1}$, for the selected radius range (varied between 40 and 130 µm for film flapping and between 30 and 65 µm for jet), and (c) the upper bound, $R_{b2}$, for the selected radius range (varied between 300 and 2000 µm for film flapping and between 1500 and 3000 µm for jet). For the lower and upper bounds, the solid line represents the distributions/set of parameters used in figure 7. Parameters are described further in table 3).
Regarding the larger drop sizes, we can also consider using a truncated Gamma distribution in the integration. In this case, the distribution of drop sizes produced by an individual bursting bubble ($p(r_{d}/\langle r_{d}\rangle,R_{b})$) would be represented by the Gamma distribution for bubbles smaller than a chosen cutoff radius $R_{b}^{max}$, and equal to zero for bubbles larger than this cutoff size. Figure 11 shows the comparison between jet drop size distributions computed using the full Gamma distribution and the truncated version, which is set to zero for bubbles larger than the chosen cutoff radius of $R_{b}^{max}$ = 2 mm. Using this truncated Gamma distribution, the predicted drop size distributions describe the data for the large drop sizes very well.
Figure 11: Measured and predicted drop size distributions ($N_{d}(r_{d})$) for the Broad-2 mm (left) and Narrow-2 mm (right) cases. The predicted jet drop size distributions are computed using the full Gamma distribution for all bubble radii (dotted line, also shown in figure 7) and the truncated Gamma distribution (solid line). The truncated Gamma distribution is equal to zero for bubbles of $R_{b}>$ 2 mm.